OpenStudy (axel.caballero):

Identify the 42nd term of an arithmetic sequence where a 1 = -12 and a 27 = 66

OpenStudy (axel.caballero):

It has to do with Arithmetic Sequences

OpenStudy (axel.caballero):

@ganeshie8 @helpme1.2 @tester97 @mathmale @destinymasha @oldrin.bataku @tester97 @starr_dynastyT @madds96 @linh412986 @chucky00 @jimboslice @figureskater120 @ryanghese12790

OpenStudy (axel.caballero):

@satellite73 @Hero @ganeshie8 @TuringTest @Nurali @ash2326 @Luigi0210 @abb0t @kelliegirl33 @eliassaab @ranga @blues @ashleyisakitty @nubeer @FutureMathProfessor @shrutipande9 @chooch146 @allie_bear22

OpenStudy (anonymous):

\[ a_n=a_1+(n-1) r\\ a_{27} =-12 + 26 r=66\\ r=3\\ a_n=-12+3(n-1) \ \] Can you find now \( a_{42}\)?

OpenStudy (anonymous):

\[ a_{42}=12+ 3(41)=-12+ 123=111 \]